# Nonlinear phase synchronization and the role of spacing in shell models

### v1

@@ II.2 Phase and Flux
which is maximized for $\varphi_{n}=\varphi_{n+1}=\pi/2$. This means that the Kolmogorov cascade with a power law spectrum and a constant energy flux is maximized if the triads are in phase at $\varphi=\pi/2$. Note that since $c$ is negative, the choice $\varphi_{n}=\pi/2$ maximizes the flux regardless of the choice of $\left|\chi_{n}\right|$ as well. This allows the system to deviate locally from the Kolmogorov solution. Similarly an inverse cascade is implied if $\varphi_{n}=\varphi_{n+1}=-\pi/2$ etc. regardless of the choice of the power law.
This means that the particular arrangement of the nonlinear dynamics of the consecutive triads is important in determining both the direction and magnitude of energy transfer through the “boundary” between two triads. In the following, we try to highlight this connection, considering different types of shell models and studying the role of triadic phases, and their distributions in the cascade process.
@@ II.2.1 Forward cascade and intermittency
In a shell model with local interactions, that conserve the inviscid invariants typical of 3D turbulence, corresponding to Eq. (1), and with the standard shell spacing $g=2$, the system is well known to drive a forward energy cascade, displaying an energy spectrum consistent with the Kolmogorov $-5/3$ scaling and a level of intermittency, somewhat consistent with experiment and direct numerical simulations [16, 15, 17].
The multifractal corrections, implied by this intermittent behavior (see, e.g., Refs. [16, 15, 41, 19]), violating Kolmogorov’s original self-similarity hypotesis, manifest themselves as non-gaussian and scale-dependent distributions of velocity fluctuations in the inertial range, with fatter tails at smaller scales. For shell models, this is undoubtedly a consequence of the time evolution, since these models lack any information about spatial distribution of singularities. Note that, rare events in time evolution of shell models are linked to sudden bursts of flux that goes through the chain of local triads that constitute them. Those events require both large amplitudes and some kind of local phase synchronization so that the burst can propagate across the chain in the correct direction.
More generally, the degree of locality of the cascade processes in shell models, can be investigated through multi-shell correlation functions [42], or via the statistical properties of the Kolmogorov multipliers, defined as $w_{n}=\frac{U_{n}}{U_{n-1}}$, and the triadic phases $\varphi_{n}$, as done in Refs. [43, 44, 45], investigating whether or not the probability distribution functions of those quantities are universal in the inertial range ( i.e. independent of the shell index). It is also worth noting that the correlation of the Kolmogorov multipliers of two different shells decays faster than that of the triadic phases [44]. This is significant, because it suggests that the long-range interactions implied by the triadic phase equation [i.e., Eq. 5] introduce nontrivial phase correlations that are particularly relevant for the statistics of extreme events causing intermittency.
@@
In Figure 1, we present the results for a standard GOY model of three dimensional Navier-Stokes, forced at large scales, or small $n$, which is extensively studied in the past [16, 15, 41]. The dynamics of the model exhibit a forward energy cascade with clear inertial range, with a spectral slope $\left|u_{n}\right|^{2}\sim k_{n}^{-2/3}$ that is the shell model equivalent of the $-5/3$ power law solution of Kolmogorov. In the inset plot the normalized PDFs of the triad phases $\varphi_{n}$ are displayed for different shells in the inertial range, showing that they collapse to a scale invariant universal distribution.
This distribution exhibits a distinct peak at the fixed point $+\pi/2$ consistent with a direct energy cascade. As we will discuss later on in Sec. III, this invariance is progressively disrupted, together with an increase in intermittency corrections, as the inter shell spacing is reduced.
@@ II.3 Sync and Kuramoto Order parameter
As discussed in Section II.1, the cascade in a shell model happens via self-organization of the triadic phases $\varphi_{n}$ of the consecutive triads that constitute it, in order to accommodate the flux that goes through them. For the forward cascade, this causes the triadic phases to cluster around $\pi/2$, and for the standard GOY model, this implies a scale-invariant universal distribution of their statistics across different shells. In the usual scenario, such statistics underlies intermittent dynamics, corresponding to rare but extreme events in the inertial range.
@@
In order to quantify the level of coherence of the chain of triads that are driving the energy cascade, we can define a global order parameter from a dynamical-systems perspective as a population of nonlinear oscillators represented by triadic phases, which would allow us to investigate the role of phase aligned triads in extreme flux events.
@@
For a 1D system of interacting oscillators, which will be described by triadic phases, we can define the Kuramoto order parameter [46, 47, 48] as:
$Re^{i\Phi}=\frac{1}{N_{t}}\sum_{t}e^{i\varphi_{t}}\;\text{,}$
where the sum $\sum_{t}$ is over the set of triads. Here, we will typically take the sum to be over the triads of the inertial range.
@@
Figure 2: Normalized probability distribution functions of the Kuramoto order parameter $R$ and $\Phi$, for various shell models exhibiting different type of cascades: a classical GOY model displaying a forward energy cascade; a helical shell model with elongated triads driving an inverse cascade (see Sec. IV and V.0.1 for a detailed discussion and the definition of the Kuramoto parameter in the context of helical shell models); and the model investigated in Ref. [49] that drives an inverse cascade via local triads (see V.0.2 below).
@@
At a given time $t$, the complex phase of the Kuramoto parameter $\Phi\left(t\right)\in[0,2\pi)$ measures the average phase of the population of triads that are considered, and the amplitude $R\left(t\right)\in[0,1]$ quantify the phase coherence of the chain. A completely incoherent state corresponds to $R(t)\approx 0$, since all the triads are uniformly distributed on the unit circle, whereas, the case with all the triad phases being perfectly aligned corresponds to $R=1$. Using the Kuramoto parameter for a set of triadic phases, was initially introduced in the context of the Burgers equation [30, 31, 50], in order to analyze the energy flux enhancement in the presence of highly coherent states, related to collisions of shocks in physical space. Here, we apply it to shell models, for characterizing the features of the phase dynamics in driving the underlying cascade and the formation of rare events.
@@
In Figure 2 we present the normalized PDF of both $R$, and $\Phi$ for the standard GOY model as an example of the forward cascade, discussed above and for two other models that are capable of generating inverse cascade (which will be discussed in more detail in Sec. V, since inverse cascade in shell models is a complicated issue). We note from the PDF of $R$ that the forward energy cascade, exhibits significantly higher level of coherence with respect to the inverse cascades in which highly coherent events are much less frequent. Regarding the normalized PDF of $\Phi$, a strong evidence of phase clustering is present for the forward cascade shell model showing a sharp distribution with a maxima at $\pi/2$, in contrast to the two models that result in inverse cascade displaying a broader, less pronounced distributions. Among the two models (that present no sign of intermittency), a weak clustering of the triadic phases around $-\pi/2$ appear in both cases. However, this is so weak that it is not clearly visible for the helical inverse cascade model in the presented plot, as compared to the model of Tom and Ray [49].
@@ II.3.1 Local synchronization and energy transfer
The clustering of triadic phases around $\pi/2$, and a significant value of the Kuramoto parameter, as in the forward cascade case shown above, provides an indicator of the global phase coherence in the entire inertial range. We can also analyze the local degree of synchronization within the triad chain and its correlation to energy flux at the corresponding shell $n$, to better understand the contribution of phase alignments in the cascade mechanism and its variations.
@@
$R_{n}e^{i\Phi_{n}}=\frac{\sum_{t\in\Gamma(n)}w_{n}e^{i\varphi_{n}}}{\sum_{\in\Gamma(n)}w_{n}}\;\text{,}$ (7)
@@
where the summation $\sum_{\in\Gamma(n)}w_{n}$ represents a -shell dependent- weighted sum. A consistent choice, in order to quantify the local coherence is to set the weight of all the triadic phases that are irrelevant to the evolution of the triadic phase $\varphi_{n}$, to zero while assigning a weight equal to the number of times a given triad phase appears on the right-hand side of the equation for the triadic phase $\varphi_{n}$ [i.e. Eq. (5)]. With this definition, a local coherent state, i.e. $R_{n}=1$, implies local phase locking, corresponding to a fixed point of the triadic phase equation. In the case of a local shell model, this definition corresponds to a sum over a list of 5 consecutive triads with the weights given by $\left\{w_{n-2},w_{n-1},w_{n},w_{n+1},w_{n+2}\right\}=\left\{1,2,3,2,1\right\}$.
@@
As shown in Fig. 3, panel (a), the PDFs of the spectral energy flux $\Pi_{n}^{E}$, computed for different shells in the inertial range, displays a typical form characterized by a pronounced peak, corresponding to a relatively small flux value (with a mean value around $0.7$ as can be seen in the inset of Figure 1), accompanied by heavy tails accounting for intense forward energy transfer events, which can reach values as large as 500 times the mean flux. This suggests the existence of rare events of sudden large transfer.
@@
The panel (b) of the same figure, shows the PDFs of the local synchronization parameter $R_{n}$, computed using Eq. (7), for different shells. Notably, the distributions are scale-independent and imply a clear tendency of the dynamics to favor locally phase-coherent configurations over incoherent ones.
@@
In panel (c), we consider the joint PDF of these two quantities for a given shell in the inertial range, in order to investigate if synchronization correlates locally with energy transfer. The joint statistics shows that while typical flux events, close to the mean value are distributed across the whole range of $R_{n}$, rare but extreme flux events, are correlated with coherent states, characterized by large values of $R_{n}$, with consecutive triads being partially-synchronized. In other words, while for small $R_{n}$ the flux events are distributed more or less symmetrically around the mean, an overwhelming majority of the large flux events, tend to have large $R_{n}$ values. In the same spirit, panel (d) shows the joint PDF between $\Phi_{n}$ and the local energy flux $\Pi_{n}^{E}$, illustrating that the extreme events are associated with a pronounced clustering of triad phases around $\pi/2.$ These results show clearly that rare events go hand in hand with local synchronization of triadic phases in shell models. In the following we will study the effect of these rare events on a well established measure of turbulence intermittency and how this behavior is affected by changing the inter-shell spacing $g$, which changes both the synchronizability of the triadic interactions and the number of its elements of the chain that these interactions constitute.
@@ III Intermittency
Extreme events, corresponding to the tails of the probability distribution of energy flux and of the fluctuations of shell amplitudes, are, as previously discussed, closely related to the level of phase coherence among a system of oscillators consisting of triads interacting along a chain. The statistical properties of these extreme events play a crucial role in the global behavior of the system, characterized by intermittency, manifesting itself in the deviation of the scaling of the structure functions, defined as $S_{p}=\left\langle\left|u_{n}\right|^{p}\right\rangle,$ from the scaling $S_{p}\propto k_{n}^{-p/3}$ predicted by Kolmogorov’s self-similarity hypothesis. Estimations of intermittency corrections in shell models can be obtained by studying the scaling behavior of the three point quantity defined as:
$\Sigma_{n,p}=\left\langle\left|Im\left(\chi_{n}\right)\right|^{p/3}\right\rangle\;\text{,}$
which eliminates period three oscillations of shell amplitudes. The scaling exponents $\xi(p)$ of the $p$-th order structure function can be estimated either by performing a linear fit in a log-log plot of $S_{p}(k_{n})$ versus $k_{n}$, or trough a more refined method such as the Extended Self-Similarity (ESS) [51].
The ESS approach, estimates the scaling exponents $\xi_{ESS}(p),$ as the power low scaling of order structure function of order $p$, with respect to the third order one, i.e. the scaling $S_{p}(k_{n})\sim S_{3}^{-\xi_{ESS}(p)}$ ($\Sigma_{p}(k_{n})\sim\Sigma_{3}^{-\xi_{ESS}(p)}$). This approach can reveal hidden scalings in turbulent structure functions (see Ref. [52] as an example in shell model context). The improved accuracy of the ESS method has been explained as its ability to suppress subdominant corrections Ref . [53] and an interpretation of ESS within the multifractal formalism has been proposed in Ref. [54].
In the following, we will present different scaling exponents $\xi(p)$ that are extracted form numerical simulations of different shell models, using the ESS procedure, restricting the fits to a range corresponding to the constant flux window in the inertial range. No hidden scaling regime is revealed by the ESS analysis (i.e. $\xi_{ESS}(p)\simeq\xi(p)$); instead, here it helps refine our estimation of the intermittency.
Figure 4: Scaling of the structure functions $S_{n,p}$ (orders $p=2-6$) as a function of the shell wave-number $k_{n}$, shown for two representative cases. The right panel ($g=2$) presents the usual intermittency corrections, while the left panel ($g=1.4$) presents an enhancement of intermittency. The insets display the PDFs of the triadic phases $\varphi_{n}$ for different shells in the inertial range, with colors indicating shell index $n$ as shown in the colorbar.

### v2

@@ II.2 The Role of Triadic Phase on Energy Flux
which is maximized for $\varphi_{n}=\varphi_{n+1}=\pi/2$. This means that the Kolmogorov cascade with the correspondingly power law spectrum and a constant energy flux is maximized if the triads are aligned at $\varphi_{n}=\varphi_{n+1}=\pi/2$. Note that since $c$ is negative, this choice maximizes the forward flux regardless of the choice of $\left|\chi_{n}\right|$ as well. This allows the system to deviate locally from the Kolmogorov solution. Similarly an inverse cascade is implied if $\varphi_{n}=\varphi_{n+1}=-\pi/2$ regardless of the choice of the power law.
This means that the particular arrangement of the nonlinear dynamics of the consecutive triads is important in determining both the direction and magnitude of energy transfer through the “boundary” between the two triads. In the following, we try to highlight this connection, considering different types of shell models and studying the role of triadic phases, and their statistical distributions in the cascade process.
@@ II.2.1 Forward cascade and intermittency
In a shell model with local interactions, that conserves the inviscid invariants typical of 3D turbulence, corresponding to Eq. (1), and with the standard shell spacing $g=2$, one gets a forward energy cascade, and an energy spectrum consistent with the Kolmogorov 5/3-5/3 scaling together with a level of intermittency, somewhat consistent with experiment and direct numerical simulations [16, 15, 17].
The multifractal corrections, implied by this intermittent behavior (see, e.g., Refs. 16, 15, 42, 19), violating Kolmogorov’s original self-similarity hypotesis, manifest themselves as non-gaussian and scale-dependent distributions of velocity fluctuations in the inertial range, with fatter tails at smaller scales. For shell models, this is undoubtedly a consequence of the time evolution, since these models lack any information about spatial distribution of singularities. Note that, rare events in the time evolution of shell models are linked to sudden bursts of flux that goes through the chain of local triads that constitute them. Those events require both large amplitudes and some kind of local phase synchronization so that the burst can propagate across the chain in the correct direction.
More generally, the degree of locality of the cascade processes in shell models, can be investigated through multi-shell correlation functions [43], or via the statistical properties of the Kolmogorov multipliers, defined as $w_{n}=\frac{U_{n}}{U_{n-1}}$, and the triadic phases $\varphi_{n}$, as done in Refs. 44, 45, 46, investigating whether or not the probability distribution functions of those quantities are universal in the inertial range ( i.e. independent of the shell index). It is also worth noting that the correlation of the Kolmogorov multipliers of two different shells decays faster than that of the triadic phases [45]. This is significant, because it suggests that the longer range interactions implied by the triadic phase equation [i.e., Eq. (5)] introduce nontrivial phase correlations that are particularly relevant for the statistics of extreme events causing intermittency.
@@
In Figure 1, we present the results for a standard GOY model of three dimensional Navier-Stokes turbulence, forced at large scales, or small $n$, which is extensively studied in the past [16, 15, 42]. The dynamics of the model exhibit a forward energy cascade with clear inertial range, with a spectral slope $\left|u_{n}\right|^{2}\sim k_{n}^{-2/3}$, which is the shell model equivalent of $E(k)\propto k^{-5/3}$. In the inset plot, the normalized PDFs of the triad phases $\varphi_{n}$ are displayed for different shells in the inertial range, showing that they collapse to a more or less scale invariant universal distribution similarly to Kolmogorov multipliers [45].
This distribution exhibits a distinct peak at the fixed point $\pi/2$ consistent with a direct energy cascade. As we will discuss later in Sec. III, reducing the inter-shell spacing progressively disrupts scale invariance and enhances intermittency corrections.
@@ II.3 Sync and Kuramoto Order parameter
As discussed in Section II.1, the cascade in a shell model happens via self-organization of the triadic phases $\varphi_{n}$ of the consecutive triads that constitute it, in order to accommodate the flux that goes through them. For the forward cascade, this causes the triadic phases to cluster around $\pi/2$, and for the standard GOY model, this implies a scale-invariant universal distribution of their statistics in the inertial range. In the usual scenario, such statistics underlies intermittent dynamics, corresponding to rare but extreme events.
@@
In order to quantify the level of coherence of the chain of triads that facilitates the energy cascade, we can define a global order parameter by considering the population of nonlinear oscillators represented by triadic phases, in order to explore the role of phase aligned triads in extreme flux events.
@@
For a 1D system of interacting nonlinear oscillators, which are described by triadic phases, we can define the Kuramoto order parameter [47, 48, 49] as:
$Re^{i\Phi}=\frac{1}{N_{m}}\sum_{m}e^{i\varphi_{m}}\;\text{,}$ (7)
where the sum $\sum_{m}$ is over all the triads in the inertial range, where a clear power low scaling is observed, independent of forcing and dissipation.
@@
Figure 2: Normalized probability distribution functions of the Kuramoto order parameter $R$ and $\Phi$, for various shell models exhibiting different type of cascades: a classical GOY model displaying a forward energy cascade; a helical shell model with elongated triads driving an inverse cascade (see Sec. IV and V.0.1 for a detailed discussion and the definition of the Kuramoto parameter in the context of helical shell models); and the model investigated in Ref. 50 that drives an inverse cascade via local triads (see V.0.2 below).
@@
At a given time $t$, the complex phase of the Kuramoto parameter $\Phi\left(t\right)\in[0,2\pi)$ measures the average phase of the population of triads that are considered, and the amplitude $R\left(t\right)\in[0,1]$ quantify the phase coherence of the chain. A completely incoherent state corresponds to $R(t)\approx 0$, since all the triads are uniformly distributed on the unit circle, whereas, the case with all the triad phases being perfectly aligned corresponds to $R=1$. Using the Kuramoto parameter for a set of triadic phases, was initially introduced in the context of the Burgers equation [30, 31, 51] in order to analyze the energy flux enhancement due to highly coherent states, related to collisions of shocks in physical space. Here, we apply it to shell models, for characterizing the dynamics of triadic phases in driving the underlying cascade and the formation of rare events that cause intermittency.
@@
In Figure 2 we present the normalized PDF of both $R$, and $\Phi$ for the standard GOY model that drives a forward cascade, discussed above and for two other models that are capable of generating inverse cascade (which will be discussed in more detail in Sec. V, since inverse cascade in shell models is a complicated issue). Note that the PDF of $R$ for the forward cascade peaks at around $R\approx 0.7$, it exhibits significantly higher level of coherence with respect to the two inverse cascade models, which peak at $R\approx 0.1$ and at $R\approx 0.25$ respectively. Regarding the normalized PDF of $\Phi$, a strong evidence of phase clustering is present for the forward cascade shell model showing a sharp distribution with a maxima at $\pi/2$, in contrast to the two models that result in inverse cascade displaying a broader, less pronounced distributions. Nevertheless, the two non-intermittent models display a weak clustering of the triadic phases around $\pi/2$. However, this is so weak that it is not clearly visible for the helical inverse cascade model in the presented plot, as compared to the model of Tom and Ray [50].
@@ II.3.1 Local synchronization and energy transfer
The clustering of triadic phases around $\pi/2$, and a significant value of the Kuramoto parameter, as in the forward cascade case shown in Fig. 2, provides an indicator of the global phase coherence in the entire inertial range. We can also analyze the local degree of synchronization within the triad chain and its correlation to energy flux at the corresponding shell $n$, to better understand the contribution of phase alignments in the cascade mechanism and its variations.
@@
$R_{n}e^{i\Phi_{n}}=\frac{\sum_{t\in\Gamma(n)}w_{t}e^{i\varphi_{t}}}{\sum_{t\in\Gamma(n)}w_{t}}\;\text{,}$ (8)
@@
where the summation $\sum_{t\in\Gamma(n)}w_{t}$ represents a shell-dependent weighted sum over a set of triads $\Gamma(n)$. A consistent choice, in order to quantify the local coherence is to restrict the summation to the set of interacting triads $\Gamma(n)$ which are connected to the central triad of the $n$-th shell. In this approach the weight of each triadic phase is equal to the number of times it appears on the right-hand side of the evolution equation for $\varphi_{n}$ [i.e., Eq. (5)]. While the weights of those who do not appear are set to zero. With this definition, a local coherent state ($R_{n}=1$), corresponds to local phase locking, which implies a fixed point of the triadic phase evolution equation. In the case of a local shell model, this definition corresponds to a sum over a list of 5 consecutive triads with the weights given by $\left\{w_{n-2},w_{n-1},w_{n},w_{n+1},w_{n+2}\right\}\equiv\left\{1,2,3,2,1\right\}$.
@@
As shown in Fig. 3, panel (a), the PDFs of the spectral energy flux $\Pi_{n}^{E}$, computed for different shells in the inertial range, display a typical form characterized by a pronounced peak, corresponding to a relatively small flux value (with a mean value around $0.7$ as can be seen in the inset of Figure 1), accompanied by heavy tails accounting for intense forward energy transfer events, which can reach values as large as 500 times the mean flux. This means that there are rare events that drive sudden and large energy transfer.
@@
The panel (b) of the same figure, shows the PDFs of the local synchronization parameter $R_{n}$, computed using Eq. (8), for different shells. Notably, the distributions are scale-independent in the inertial range and imply a clear tendency of the dynamics to favor locally phase-coherent configurations over incoherent ones.
@@
In panel (c), we consider the joint PDF of these two quantities for a given shell in the inertial range, in order to investigate if local synchronization correlates with energy transfer. The joint distribution shows that while typical flux events, close to the mean value are distributed across the whole range of $R_{n}$, rare but extreme flux events only occur for large values of $R_{n}$, where consecutive triads are partially-synchronized. In other words, while for small $R_{n}$ the flux events are distributed more or less symmetrically around the mean, an overwhelming majority of the large flux events, tend to have large $R_{n}$ values. In the same spirit, panel (d) shows the joint PDF between $\Phi_{n}$ and the local energy flux $\Pi_{n}^{E}$, illustrating that the extreme events are associated with a pronounced clustering of triad phases around $\pi/2.$ These results show clearly that rare events go hand in hand with local synchronization of triadic phases in shell models. In the following we will study the effect of these rare events on a well-established measure of turbulence intermittency and how this behavior varies with the inter-shell spacing $g$, which changes both the synchronizability of the triadic interactions and the number of its elements.
@@ III Intermittency
We have shown that extreme events in energy flux (and amplitudes $U_{n}$) are linked to triadic phase coherence. The statistical properties of these events determine the level of intermittency, manifesting itself in the deviation of the scaling of the structure functions, defined as $S_{p}(k_{n})=\left\langle\left|u_{n}\right|^{p}\right\rangle,$ from the scaling $S_{p}\propto k_{n}^{-p/3}$ predicted by Kolmogorov’s self-similarity hypothesis. In the GOY shell model, intermittency corrections can be more accurately quantified by studying the scaling behavior of the three point correlations:
$\Sigma_{p}(k_{n})=\left\langle\left|Im\left(\chi_{n}\right)\right|^{p/3}\right\rangle\;\text{,}$
which eliminates period three oscillations of shell amplitudes [15]. The scaling exponents $\xi(p)$ of the $p$-th order structure function can be estimated either by performing a linear fit in a log-log plot of $S_{p}(k_{n})$ or $\Sigma_{p}(k_{n})$ versus $k_{n}$, or trough a more refined method such as the Extended Self-Similarity (ESS) [52].
The ESS approach, estimates the scaling exponents $\xi_{ESS}(p),$ as the power low scaling of the structure function of order $p$, with respect to the third order one, i.e. the scaling $S_{p}(k_{n})\sim S_{3}^{-\xi_{ESS}(p)}$ ($\Sigma_{p}(k_{n})\sim\Sigma_{3}^{-\xi_{ESS}(p)}$). This approach can reveal hidden scalings in turbulent structure functions (see Ref. 53 as an example in the shell model context). The improved accuracy of the ESS method has been explained as its ability to suppress subdominant corrections [54] and an interpretation of ESS within the multifractal formalism has been proposed in Ref. 55.
In the following analysis, we present different scaling exponents $\xi(p)$ that are extracted from numerical simulations of different shell models. In particular, for the GOY model we employ the three-point correlation $\Sigma_{p}(k_{n})$, using the ESS procedure, restricting the fits to a range of scales corresponding to the constant flux window in the inertial range. Note that even though no hidden scaling regime is revealed by the ESS analysis (i.e. $\xi_{ESS}(p)\simeq\xi(p)$); here, it helps refine our estimation of the intermittency.
Figure 4: Scaling of the structure functions $S_{p}(k_{n})$ (orders $p=2,3,\dots 6,7$) as a function of the shell wave-number $k_{n}$, shown for two representative values of the inter-shell spacing $g=1.4$ and $g=2$. The right panel ($g=2$) presents the typical intermittency corrections, while the left panel ($g=1.4$) presents an enhancement of intermittency. The insets display the PDFs of the triadic phases $\varphi_{n}$ for different shells in the inertial range, with colors indicating shell index $n$ as shown in the colorbar.